# Federated Learning with Dynamic Client Arrival and Departure: Convergence and Rapid Adaptation via Initial Model Construction

### v2

@@ 4 Theoretical Analysis
Local loss function $F_{k}^{(g)}$ is $\beta$-smooth, $\forall k\in\mathcal{K}^{(g)},g$: $\|\nabla F_{k}^{(g)}(\bm{w})-\nabla F_{k}^{(g)}(\bm{w}^{\prime})\|\leq\beta\|\bm{w}-\bm{w}^{\prime}\|,~{}\forall\bm{w},\bm{w}^{\prime}\in\mathbb{R}^{M},$ which implies the $\beta$-smoothness of the global loss function $F^{(g)},\;\forall g\in\mathcal{G}$.
@@
There exist finite constants $\zeta_{1}\geq 1$, $\zeta_{2}\geq 0$, for any set of coefficients $\{a_{k}\geq 0\}_{k\in\mathcal{K}^{(g)}}$, where $\sum_{k\in\mathcal{K}^{(g)}}a_{k}=1$, such that $\forall g,\bm{w}$ $\sum\nolimits_{k\in\mathcal{K}^{(g)}}a_{k}\|\nabla F_{k}^{(g)}(\bm{w})\|^{2}\leq\zeta_{1}\|\sum\nolimits_{k\in\mathcal{K}^{(g)}}a_{k}\nabla F_{k}^{(g)}(\bm{w})\|^{2}+\zeta_{2}$.
The parameters $\zeta_{1}$ and $\zeta_{2}$, introduced in Assumption 2, quantify the degree of heterogeneity (i.e., the extent of non-i.i.d. behavior) among the datasets distributed across devices. When the datasets are fully homogeneous (i.e., i.i.d. across devices), these parameters take the values $\zeta_{1}=1$ and $\zeta_{2}=0$. However, as the heterogeneity of the data increases, both $\zeta_{1}$ and $\zeta_{2}$ will grow accordingly. These parameters play crucial roles in our convergence analysis, as they influence the allowable step size for SGD and the overall convergence rate of ML model training.
To further capture the heterogeneity within the local datasets of individual devices, we introduce the concept of local data variability.
@@
The local data variability at each device $k$ is measured via $\Theta_{k}\geq 0$, which $\forall\bm{w},k$ satisfies $\|\nabla f(\bm{w},d)-\nabla f(\bm{w},d^{\prime})\|\leq\Theta_{k}^{(g)}\|d-d^{\prime}\|,\quad\forall d,d^{\prime}\in\mathcal{D}_{k}^{(g)}$. We further define $\Theta^{(g)}=\max_{k\in\mathcal{K}^{(g)}}\{\Theta_{k}^{(g)}\}$.
@@
Suppose each client performs local SGD updates and the server aggregates via weighted averaging. Define $e_{\text{max}}^{(g)}=\max_{k\in\mathcal{K}^{(g)}}e_{k}^{(g)}$ and $e_{\text{min}}^{(g)}=\min_{k\in\mathcal{K}^{(g)}}e_{k}^{(g)}$. Assuming the learning rate $\eta^{(g)}$ is less than
$\min\!\Bigl{\{}(0.5-((e_{\text{min}}^{(g)})^{-1}-1)^{2})\,\beta^{-1},\;\sqrt{\Lambda^{(g)}}\,(2\beta\sqrt{(\zeta_{1}+\Lambda^{(g)})\,e_{\text{max}}^{(g)}(e_{\text{max}}^{(g)}-1)})^{-1}\Bigr{\}},$
where $\Lambda^{(g)}<1$ is a constant, then for any pattern of client arrivals and departures, the cumulative average of the gradient of the global loss over $G$ rounds satisfies:
$\frac{1}{G}\sum_{g=1}^{G}\mathbb{E}\|\nabla F^{(g)}(\bm{w}^{(g)})\|^{2}\leq\underbrace{\frac{1}{G}\sum_{g=1}^{G}\frac{2\mathbb{E}\bigl{[}F^{(g)}(\bm{w}^{(g)})-F^{(g)}(\bm{w}^{(g+1)})\bigr{]}}{\eta^{(g)}(1-\Lambda^{(g)})}}_{\text{(a)}}+\underbrace{\frac{1}{G}\sum_{g=1}^{G}\frac{4\beta(\Theta^{(g)})^{2}\eta^{(g)}}{1-\Lambda^{(g)}}\sum_{k\in\mathcal{K}^{(g)}}\frac{(\widetilde{\sigma}_{k}^{(g)})^{2}\,e_{k}^{(g)}}{B_{k}^{(g)}}\frac{D_{k}^{(g)}-B_{k}^{(g)}}{(D_{k}^{(g)})^{2}}}_{\text{(b)}}$
$\quad+\underbrace{\frac{1}{G}\Bigl{(}\sum_{g=1}^{G}\frac{4\beta^{2}(\eta^{(g)})^{2}}{D^{(g)}(1-\Lambda^{(g)})}\Bigr{)}\sum_{k\in\mathcal{K}^{(g)}}(e_{k}^{(g)}-1)\Bigl{(}1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\Bigr{)}\frac{\Theta^{(g)}(\widetilde{\sigma}_{k}^{(g)})^{2}}{B_{k}^{(g)}}}_{(c)}+\underbrace{\frac{1}{G}\sum_{g=1}^{G}8\beta^{2}(\eta^{(g)})^{2}(e_{\max}^{(g)})(e_{\max}^{(g)}-1)\frac{\zeta_{2}}{1-\Lambda^{(g)}}}_{\text{(d)}}.$
@@
Please see Appendix E. ∎
@@
Term (a) reflects the expected improvement in the loss function at each training stage. Terms (b) and (c) capture the negative impacts stemming from mini-batch sampling noise (controlled by batch size $B_{k}^{(g)}$), local data heterogeneity across clients (measured by $\Theta^{(g)}$), and the number of local SGD steps ($e_{k}^{(g)}$). Term (d) accounts for the influence of the maximum number of local SGD steps ($e_{\text{max}}^{(g)}$) and inter-client model dissimilarity (represented by $\zeta_{2}$).
Inspiration from Theorem 1 for Initial Model Construction. Term (a) in the bound (2) decomposes neatly into two contributions:
$\mathbb{E}\bigl{[}F^{(g)}(\bm{w}^{(g)})-F^{(g)}(\bm{w}^{(g)\ast})\bigr{]}\quad\text{and}\quad\mathbb{E}\bigl{[}F^{(g)}(\bm{w}^{(g)\ast})-F^{(g)}(\bm{w}^{(g+1)})\bigr{]}.$
@@
The second of these arises from the update that follows round $g$ and is beyond our control at that stage. Our algorithmic leverage lies in the first term: by constructing an initial model $\bm{w}_{\mathrm{init}}^{(g)}$ aiming to satisfy
$\mathbb{E}\bigl{[}F^{(g)}(\bm{w}_{\mathrm{init}}^{(g)})-F^{(g)}(\bm{w}^{(g)\ast})\bigr{]}\;\leq\;\mathbb{E}\bigl{[}F^{(g)}(\bm{w}^{(g)})-F^{(g)}(\bm{w}^{(g)\ast})\bigr{]},$
@@
we directly tighten the upper bound in Theorem 1 and reduce the cumulative average gradient norm.
Algorithmic Realization via Pilot-Model Similarity Weighting. When the data distribution in round $g$ closely matches that of a previous session, reusing that session’s final model as $\bm{w}_{\mathrm{init}}^{(g)}$ naturally yields the desired reduction in Term (a). In more heterogeneous or evolving settings, however, no single past model suffices. Instead, we form
$\bm{w}_{\mathrm{init}}^{(g)}=\sum_{s<g}\alpha_{s}\,\bm{w}^{(s)},$
where the weights $\{\alpha_{s}\}$ are determined by the similarity between each prior data distribution and the current one. We estimate similarity by computing pilot-model gradients: greater alignment of gradients indicates higher relevance and thus a larger $\alpha_{s}$. This similarity-weighted ensemble initialization both accelerates convergence in stationary regimes and provides robust adaptation under distribution shift, directly minimizing the controllable component of our convergence bound. From this argument, it follows that our similarity-weighted warm-start strategy is agnostic to the choice of federated learning algorithm and can be seamlessly integrated into any aggregation scheme; we validate this generality by experiments in Section 6. To make this concrete, in the next section we describe how to construct the similarity weights $\{\alpha_{s}\}$ that underpin our warm-start strategy.
5 Dynamic Initial Model Construction for Fast Adaptation
Terminology: Algorithm 1 outlines our proposed “Dynamic Initial Model Construction” algorithm. In this context, a “session” differs from the “number of global rounds.” A new session is initiated whenever there is a change in the data distribution, such as clients joining or leaving, while the client set remains constant within a session. Each session comprises at least one global round.
Training: From 8 to 11, each client in the current set performs local model training based on either a new initial model (9) or the latest model at the server (10). Upon completion of local training, each client transmits its final local model to the server, where global aggregation is performed. It is worth noting that in Algorithm 1, the local model training and the global aggregation can be performed in any desired way (e.g., FedAvg, FedProx, SCAFFOLD, and FedACG). In 12, we only save the last global model in each session, either for pilot model computation or as components of a new initial model.

### v3

@@ III-B Theoretical Analysis
Local loss function $F_{k}^{(s)}$ is $\beta$-smooth for all $k\in\mathcal{K}^{(s)}$, $\forall\bm{w},\bm{w}^{\prime}\in\mathbb{R}^{M}$, and $s$:
$\|\nabla F_{k}^{(s)}(\bm{w})-\nabla F_{k}^{(s)}(\bm{w}^{\prime})\|\leq\beta\|\bm{w}-\bm{w}^{\prime}\|,$ (6)
which implies the $\beta$-smoothness of the global loss function $F^{(s)}$ for all $s\in\{1,\dots,S\}$.
@@
There exist finite constants $\zeta_{1}\geq 1$, $\zeta_{2}\geq 0$, such that for any set of coefficients $\{a_{k}\geq 0\}_{k\in\mathcal{K}^{(s)}}$ with $\sum_{k\in\mathcal{K}^{(s)}}a_{k}=1$, the following holds for all $s$ and $\bm{w}$:
$\sum_{k\in\mathcal{K}^{(s)}}a_{k}\|\nabla F_{k}^{(s)}(\bm{w})\|^{2}\leq\zeta_{1}\big\|\sum_{k\in\mathcal{K}^{(s)}}a_{k}\nabla F_{k}^{(s)}(\bm{w})\big\|^{2}+\zeta_{2}.$ (7)
To capture the heterogeneity within the local datasets of devices participating in a specific round, we introduce the concept of local data variability:
@@
The local data variability at each device $k$ is measured via $\Theta_{k}^{(s,t)}\geq 0$, which $\forall\bm{w},k$ and $\forall d,d^{\prime}\in\mathcal{D}_{k}^{(s)}$ satisfies
$\|\nabla f(\bm{w},d)-\nabla f(\bm{w},d^{\prime})\|\leq\Theta_{k}^{(s,t)}\|d-d^{\prime}\|.$ (8)
We further define $\Theta^{(s,t)}=\max_{k\in\mathcal{K}^{(s)}}\{\Theta_{k}^{(s,t)}\}$.
@@
Suppose each device performs local SGD updates and the server aggregates via weighted averaging. Define $e_{\text{max}}^{(s,t)}=\max_{k\in\mathcal{K}^{(s)}}e_{k}^{(s,t)}$ and $e_{\text{min}}^{(s,t)}=\min_{k\in\mathcal{K}^{(s)}}e_{k}^{(s,t)}$. Assuming the learning rate $\eta^{(s,t)}$ satisfies
$\eta^{(s,t)}<$ $\min\Bigg\{\left(0.5-\left((e_{\text{min}}^{(s,t)})^{-1}-1\right)^{2}\right)\beta^{-1},$
$\left[\sqrt{\Lambda^{(s,t)}}\cdot\left(2\beta\sqrt{(\zeta_{1}+\Lambda^{(s,t)})e_{\text{max}}^{(s,t)}(e_{\text{max}}^{(s,t)}-1)}\right)\right]^{-1}\Bigg\},$
where $\Lambda^{(s,t)}<1$ is a constant, then for any pattern of device arrivals and departures, the cumulative average of the gradient of the global loss over $S$ sessions and $T$ rounds satisfies (9).
$\hskip-42.67912pt\begin{aligned} &\frac{1}{ST}\sum_{s=1}^{S}\sum_{t=1}^{T}\mathbb{E}\|\nabla F^{(s)}(\bm{w}^{(s,t)})\|^{2}\leq\underbrace{\frac{1}{ST}\sum_{s=1}^{S}\sum_{t=1}^{T}\frac{2\mathbb{E}\big[F^{(s)}(\bm{w}^{(s,t)})-F^{(s)}(\bm{w}^{(s,t+1)})\big]}{\eta^{(s,t)}(1-\Lambda^{(s,t)})}}_{\text{(a)}}+\underbrace{\frac{1}{ST}\sum_{s=1}^{S}\sum_{t=1}^{T}\frac{4\beta(\Theta^{(s,t)})^{2}\eta^{(s,t)}}{1-\Lambda^{(s,t)}}\sum_{k\in\mathcal{K}^{(s)}}\frac{(\widetilde{\sigma}_{k}^{(s,t)})^{2}e_{k}^{(s,t)}}{B_{k}^{(s,t)}}\frac{D_{k}^{(s)}-B_{k}^{(s,t)}}{(D_{k}^{(s)})^{2}}}_{\text{(b)}}\\ &+\underbrace{\frac{1}{ST}\Bigg(\sum_{s=1}^{S}\sum_{t=1}^{T}\bigg(\frac{4\beta^{2}(\eta^{(s,t)})^{2}}{D^{(s)}(1-\Lambda^{(s,t)})}\sum_{k\in\mathcal{K}^{(s)}}(e_{k}^{(s,t)}-1)\left(1-\frac{B_{k}^{(s,t)}}{D_{k}^{(s)}}\right)\frac{\Theta^{(s,t)}(\widetilde{\sigma}_{k}^{(s,t)})^{2}}{B_{k}^{(s,t)}}\bigg)\Bigg)}_{(c)}+\underbrace{\frac{1}{ST}\sum_{s=1}^{S}\sum_{t=1}^{T}8\beta^{2}(\eta^{(s,t)})^{2}(e_{\text{max}}^{(s,t)})(e_{\text{max}}^{(s,t)}-1)\frac{\zeta_{2}}{1-\Lambda^{(s,t)}}}_{\text{(d)}}.\end{aligned}\hskip-39.83385pt$ (9)
@@
We present a sketch proof here and provide the detailed proof in our online technical report [chang2025federatedlearningdynamicclient]. The convergence analysis relies on the $\beta$-smoothness of the global loss function $F^{(s)}(\cdot)$ and the bounding of error terms introduced by stochastic sampling and local updates as follows:
Step 1. Smoothness and Error Decomposition: Let $\bm{w}^{(s,t)}$ be the global model at round $t$ of session $s$. Using the smoothness assumption, the expected descent in one global round is bounded by:
$\mathbb{E}$ $[F^{(s)}(\bm{w}^{(s,t+1)})]\leq F^{(s)}(\bm{w}^{(s,t)})$
$-\eta^{(s,t)}\mathbb{E}\left\langle\nabla F^{(s)}(\bm{w}^{(s,t)}),\bm{\Delta}^{(s,t)}\right\rangle+\frac{\beta(\eta^{(s,t)})^{2}}{2}\mathbb{E}\|\bm{\Delta}^{(s,t)}\|^{2},$ (10)
where $\bm{\Delta}^{(s,t)}=\sum_{k\in\mathcal{A}^{(s,t)}}\frac{D_{k}^{(s)}}{\sum_{j\in\mathcal{A}^{(s,t)}}D_{j}^{(s)}}\nabla\overline{F}_{k}^{(s)}$ is the aggregated update direction derived from the participating set $\mathcal{A}^{(s,t)}$. Using the algebraic identity $2\langle\bm{a},\bm{b}\rangle=\|\bm{a}\|^{2}+\|\bm{b}\|^{2}-\|\bm{a}-\bm{b}\|^{2}$, we decompose the inner product to isolate the “true” gradient norm. This yields:
$\mathbb{E}[F^{(s)}(\bm{w}^{(s,t+1)})]\leq F^{(s)}(\bm{w}^{(s,t)})-\frac{\eta^{(s,t)}}{2}\|\nabla F^{(s)}(\bm{w}^{(s,t)})\|^{2}$
$+\underbrace{\frac{\beta(\eta^{(s,t)})^{2}}{2}\mathbb{E}\|\bm{\Delta}^{(s,t)}\|^{2}}_{\text{(A) Gradient Variance}}+\underbrace{\frac{\eta^{(s,t)}}{2}\mathbb{E}\|\nabla F^{(s)}(\bm{w}^{(s,t)})-\bm{\Delta}^{(s,t)}\|^{2}}_{\text{(B) Model Drift}}.$ (11)
Equation (III-B) highlights the two forces opposing convergence: the variance of the stochastic gradients (Term A) and the drift between the local accumulated gradients and the true global gradient (Term B).
Step 2. Bounding Gradient Variance (Term A): Unlike standard SGD analysis, our sampling is performed without replacement. Utilizing the properties of the hypergeometric distribution (finite population correction), we bound the variance of the mini-batch gradients. Specifically, we show that Term A has an upper bound:
$\beta(\eta^{(s,t)})^{2}(\Theta^{(s,t)})^{2}\sum_{k\in\mathcal{K}^{(s)}}\frac{(\widetilde{\sigma}_{k}^{(s,t)})^{2}}{D_{k}^{(s)}}\sum_{e=1}^{e_{k}^{(s,t)}}\frac{1-\frac{B_{k}^{(s,t)}}{D_{k}^{(s)}}}{B_{k}^{(s,t)}}.$ (12)
Crucially, the factor $(1-B_{k}^{(s,t)}/D_{k}^{(s)})$ captures the variance reduction: as the local batch size $B_{k}^{(s,t)}$ approaches the local dataset size $D_{k}^{(s)}$, the variance contribution vanishes.
Step 3. Bounding Model Drift (Term B): Since devices perform $e_{k}^{(s,t)}$ local steps, the local model $\bm{w}_{k}^{(s,t),e}$ drifts from $\bm{w}^{(s,t)}$. We bound this drift by recursively summing the squared norms of differences across local steps. The accumulated drift is bounded by:
$\text{Term (B)}\leq\eta^{(s,t)}\mathcal{E}_{\text{var}}+\eta^{(s,t)}\mathcal{E}_{\text{grad}}\|\nabla F^{(s)}(\bm{w}^{(s,t)})\|^{2},$ (13)
where $\mathcal{E}_{\text{var}}$ represents drift caused by noise and $\mathcal{E}_{\text{grad}}$ represents drift caused by the gradient direction itself.
Step 4. Final Convergence Bound: Substituting the bounds for Terms (A) and (B) back into (III-B), we group the coefficients of $\|\nabla F^{(s)}(\bm{w}^{(s,t)})\|^{2}$. To guarantee convergence, we require the step size $\eta^{(s,t)}$ to be sufficiently small such that the net coefficient is negative:
$\eta^{(s,t)}\leq\frac{\sqrt{\Lambda^{(s,t)}}}{2\beta\sqrt{\left(\zeta_{1}+\Lambda^{(s,t)}\right)e_{\text{max}}^{(s,t)}\left(e_{\text{max}}^{(s,t)}-1\right)}},$ (14)
where $\Lambda^{(s,t)}<1$ is a control parameter. Under this condition, telescoping the sum over $S$ sessions and $T$ rounds and rearranging yields the bound on $\frac{1}{ST}\sum_{s=1}^{S}\sum_{t=1}^{T}\|\nabla F^{(s)}(\bm{w}^{(s,t)})\|^{2}$ in Theorem 1. ∎
@@
Examining (9), term (a) reflects the expected improvement in the loss function at each training stage. (b) and (c) capture the negative impacts stemming from mini-batch sampling noise (controlled by batch size $B_{k}^{(s,t)}$), local data heterogeneity across devices (measured by $\Theta^{(s,t)}$), and the number of local SGD steps ($e_{k}^{(s,t)}$). Term (d) accounts for the influence of the maximum number of local SGD steps ($e_{\text{max}}^{(s,t)}$) and inter-device model dissimilarity (represented by $\zeta_{2}$).
@@
Implications: Term (a) in (9) decomposes neatly into two contributions: $\mathbb{E}\bigl[F^{(s)}(\bm{w}^{(s,t)})-F^{(s)}(\bm{w}^{(s)\ast})\bigr]$ and $\mathbb{E}\bigl[F^{(s)}(\bm{w}^{(s)\ast})-F^{(s)}(\bm{w}^{(s,t+1)})\bigr].$ The second term arises from the update following round $t$ and is beyond our control at that stage. Our algorithmic leverage lies in the first term, specifically at the session boundary ($t=1$). By constructing an initial model $\bm{w}_{\text{init}}^{(s,1)}$ aiming to satisfy
$\mathbb{E}\bigl[F^{(s)}(\bm{w}_{\text{init}}^{(s,1)})-F^{(s)}(\bm{w}^{(s)\ast})\bigr]$ (15)
$\leq\ \mathbb{E}\bigl[F^{(s)}(\bm{w}^{(s-1,T+1)})-F^{(s)}(\bm{w}^{(s)\ast})\bigr],$
@@
we directly tighten the upper bound in Theorem 1 and reduce the cumulative average gradient norm. We next develop our algorithm for constructing the initial model $\bm{w}_{\text{init}}^{(s,1)}$ effectively.
IV Dynamic Initial Model Construction